OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 3m - 5 = 7m - 21 Solve. One point for showing the steps to solve each equation and one point for the correct solution. 3m - 5 = 7m - 21 @Mathematics

jimthompson5910 (jim_thompson5910):

3m-5=7m-21 3m=7m-21+5 3m-7m=-21+5 -4m=-21+5 -4m=-16 m=-16/(-4) m=4